OpenStudy (anonymous):

What are the possible values of x in 8x2 + 4x = -1?

OpenStudy (anonymous):

What part are you having troubles with, human?

OpenStudy (anonymous):

*calculates* You will get complex solutions

OpenStudy (anonymous):

\[4x(2x+1)=-1\] \[x(2x+1) = -\frac{ 1 }{ 4 }\] \[2x^2+x = -\frac{ 1 }{ 4 }\] \[x^2+\frac{ x }{ 2 }+\frac{ 1 }{ 16 }=-\frac{ 1 }{ 16 }\] \[\left( x+\frac{ 1 }{ 4 } \right)^2=-\frac{ 1 }{ 16 }\]

OpenStudy (anonymous):

\[x= -\frac{ 1 }{ 4 }+\frac{ i }{ 4 }\]